OpenStudy (anonymous):

Solve the rational equation. Express numbers as integers or simplified fractions. 2/k^2-9k+14 - 6/k-7=1/4k-8

jhonyy9 (jhonyy9):

k^2 -9k +14 =(k-2)(k-7) 4k-8=4(k-2) so find the common denominator and multiplie the numerators with necessary factors make the calcules and will get the right answer sure easy good luck hope so much that is understandably what i have wrote above like a little help

OpenStudy (anonymous):

Please help. This is my last question on the section.

hero (hero):

Would you mind explaining exactly what you need help with regarding this problem?

OpenStudy (anonymous):

The whole problem.

hero (hero):

I see. I can show you how to do it and explain the steps, but on a different site.

OpenStudy (anonymous):

What is the site?

hero (hero):

Hang on a minute. I will send you a link.